OpenStudy (anonymous):

A funnel is shaped like a cone and is 4.5 inches high and has a diameter of 6 inches. What is the volume of the funnel? Use 3.14 for pi. Simplify to the nearest hundredth.

OpenStudy (egbeach):

whats the formula for finding the volume of a cone?

OpenStudy (egbeach):

\[\Pi r ^{2}\frac{ h }{ 3 }\] \[3^{2}\frac{ 4.5 }{ 3 }\]

OpenStudy (egbeach):

9x1.5pi =13.5pi= about 42.41

OpenStudy (astrophysics):

 notice the radius is half of the diameter so you have \[r = \frac{ d }{ 2 } = \frac{ 6 }{ 2 } = 3\] \[V = (3.14)(3)^2\left( \frac{ 4.5 }{ 3 } \right) inches^3\]